### v2

# Non-Markovian Multiphoton Chiral Dynamics with Giant Systems

Abstract
Chiral interactions enable directional control of quantum light, providing new routes for nonreciprocal photon dynamics. While previous studies focused on single photon regimes, this work explores the non-Markovian multiphoton regime in a giant nonlinear system nonlocally coupled to a one-dimensional waveguide, with a tunable phase difference breaking parity symmetry. We show that single-photon transport remains reciprocal, but multiphoton dynamics become chiral, generating direction-dependent higher-order correlations and enabling deterministic creation of multiphoton states with nontrivial statistics from a single input direction, even under strong dissipation. The non-Markovian nature allows tuning of transmitted photon statistics via the coupling-point separation. At a specific phase, internal dynamics become Markovian while output photons retain non-Markovian features. Under strong coherent driving, we uncover nonreciprocal dissipative phase transitions and demonstrate how non-Markovianity shapes the critical response. This framework offers a general strategy for controlling chiral multiphoton dynamics in nonlocally coupled waveguide QED systems.
I Introduction
The ability to engineer single-photon-level light-matter interactions in confined geometries is essential for quantum information processing, with applications ranging from cat-state generation to quantum communication protocols Milonni (2013); Kimble (1998); Walls and Milburn (2008); Haroche and Raimond (2006). A particularly promising direction is chiral quantum optics, where emitters couple directionally to propagating modes in one-dimensional (1D) media Gardiner (1993); Carmichael (1993); Soro and Kockum (2022); Lodahl et al. (2017); Roushan et al. (2017); Kannan et al. (2023). Such direction-dependent interactions offer new routes for generating exotic quantum many-body states Stannigel et al. (2012); Ramos et al. (2014) and realizing cascaded quantum networks Cirac et al. (1997); Kimble (2008). Early realizations of chiral coupling relied on nanophotonic systems, where polarization-selective waveguides enforced directionality at the single-emitter level Petersen et al. (2014); Sayrin et al. (2015); Scheucher et al. (2016); Lodahl et al. (2017). In these systems, transitions between certain emitter states can be induced exclusively by light traveling in a specific direction.
A further promising route to achieving chiral interactions is through a waveguide QED architecture coupled to a giant emitter Frisk Kockum et al. (2014); Guo et al. (2017); Kockum et al. (2018); Kannan et al. (2020); Soro and Kockum (2022); Cheng et al. (2022); Noachtar et al. (2022); Terradas-Briansó et al. (2022); Du et al. (2021); Joshi et al. (2023). Unlike “small” emitters, giant emitters couple to multiple, spatially separated points along a waveguide. Owing to interference between propagating fields between the coupling points, giant emitters can exhibit a rich variety of physical phenomena, including a frequency-dependent Lamb shift Frisk Kockum et al. (2014), nonexponential atomic decay and distinctive single- and two-photon scattering properties Guo et al. (2017), topological and retardation effects Cheng et al. (2022), and nonreciprocal frequency conversion Du et al. (2021). Experimentally, recent advances in superconducting platforms have enabled the realization of giant artificial atoms based on transmon qubits Gu et al. (2017); Kannan et al. (2020); Joshi et al. (2023). It has been shown that time-modulated interactions allow precise control over the relative phase $\phi$ between coupling points Joshi et al. (2023). This nonzero $\phi$ breaks parity symmetry and enables tunable, chiral light–matter interactions.
Despite this parity-symmetry breaking, single-photon transport remains reciprocal due to an underlying parity-time ($\mathcal{PT}$) symmetry Guimond et al. (2020); Wang et al. (2021); Chen et al. (2022); Cheng et al. (2022); Wang and Li (2022); Roccati and Cilluffo (2024). That is, the reflection amplitude is identical for left- and right-incident single photons. In contrast, multi-photon processes break this reciprocity. In this work, we explore the nonlinear, driven dynamics of a giant chiral system in the few- and many-photon regimes, fully incorporating both nonlinearity and retardation effects Shi et al. (2015); Gu et al. (2024). For nonzero $\phi$, input photons from one direction can experience complete transparency—propagating coherently without interacting with the system Joshi et al. (2023). Conversely, when driven from the opposite direction, we find that the same system generates output with strongly nonclassical photon statistics.
In the weak-driving regime, the emergence of non-Poissonian photon statistics is governed by the system’s intrinsic nonlinearity, while their deterministic generation is protected by $\mathcal{PT}$ symmetry. Surprisingly, even when the nonlinear scale is smaller than the decay rate—deep in the dissipative regime—quantum interference within the nonlocal coupling region leads to highly nontrivial correlations Chang et al. (2016); Kong et al. (2025). These correlations are direction-dependent Yang and Chang (2024) and sensitive to the spatial separation of the coupling points, reflecting the inherent non-Markovianity of the system. Specifically, at $\phi=\pm\mathrm{\pi}/2$ Roccati and Cilluffo (2024), interference between coupling points vanishes, and the system’s dynamics become exactly Markovian. This Markovian behavior is not limited to the single-excitation case Roccati and Cilluffo (2024), but persists regardless of the excitation number. In this limit, a master equation description becomes exact Gardiner and Zoller (2004); Walls and Milburn (2008), allowing for full characterization under arbitrary driving. Yet, the output fields remain non-Markovian due to their interaction with the extended coupling structure. This interplay between internal Markovianity and external memory effects enables controlled realization of nonreciprocal dissipative phase transitions (DPTs) Houck et al. (2012); Kessler et al. (2012); Diehl et al. (2008); Carmichael (2015); Fitzpatrick et al. (2017); Aspelmeyer et al. (2014); Benito et al. (2016); Minganti et al. (2018). By tuning the chiral phase and leveraging nonlocal interference, our framework establishes a versatile platform for engineering nonreciprocal, non-Markovian multiphoton dynamics in waveguide QED.
II Results
As a paradigmatic example, we use a giant Kerr cavity Imamoḡlu et al. (1997); Rebić et al. (2009); Blais et al. (2021), which reduces to a giant two-level atom as its nonlinearity $U$ approaches infinity, to demonstrate the multiphoton chiral dynamics. But it can also be extended to other giant systems. The setup, illustrated in Fig. 1, involves a giant Kerr cavity nonlocally coupled to a 1D bath (the waveguide) at two points, 0 and $d$. Without loss of generality, we introduce a phase shift $\varphi$ at coupling point 0, as only the phase difference between the two coupling points is relevant. The full Hamiltonian $H=H_{0}+H_{1}$, where
$H_{0}=\frac{U}{2}b^{{\dagger}2}b^{2}+H_{\lambda\mathrm{D}}+\sum_{\lambda=\mathrm{L,R}}\int_{-\infty}^{+\infty}\varepsilon_{\lambda k}a_{\lambda k}^{{\dagger}}a_{\lambda k}\mathrm{d}k,$ (1)
is the free Hamiltonian of the cavity and the waveguide respectively, plus the driving term $H_{\lambda\mathrm{D}}=\left(\Omega_{\lambda}\left(k_{i}\right)b^{\dagger}e^{-i\varepsilon_{\lambda k_{i}}t}+H.c.\right)$, where $b$ ($a_{\lambda k}$) is the annihilation operator for the cavity mode ($\lambda$-moving bath mode with momentum $k$), and their interaction is depicted by
$H_{1}=\sum_{\lambda=\mathrm{L,R}}\int_{-\infty}^{+\infty}V_{\lambda k}a_{\lambda k}b^{{\dagger}}\mathrm{d}k+\mathrm{H.c..}$ (2)
Here, $\Omega_{\lambda}\left(k_{i}\right)=\Omega_{0}\eta_{\lambda k_{i}}\left(\phi\right)$ is the Rabi frequency of the driving field propagating in direction $\lambda$ with momentum $k_{i}$, where $\Omega_{0}$ is a constant strength, and $\eta_{\lambda k_{i}}\left(\phi\right)=e^{i\phi}+e^{i\theta_{\lambda k_{i}}}$, with the propagation phase $\theta_{\lambda k_{i}}=\left(\sigma_{\lambda}k_{0}+k_{i}\right)d$, $\sigma_{R}=1$, and $\sigma_{L}=-1$. The Hamiltonian is written in a rotating frame with respect to the central frequency $k_{0}$ of the bath, which is assumed to coincide with the cavity frequency for simplicity. The $k$-dependent coupling strength $V_{\lambda k}=V_{0}\eta_{\lambda k}\left(\phi\right)$, where $V_{0}$ is a constant strength. The group velocity is taken to be unity. The spectrum of the bath is assumed to be linear, i.e., $\varepsilon_{\lambda k}=\sigma_{\lambda}k$, but it can be extended to other spectra, such as cosine spectrum Chang et al. (2011); Shi et al. (2016) and Ohmic spectrum Shi et al. (2018).
Figure 1: Schematic of a giant Kerr cavity coupled to a waveguide at two coupling points separated by a distance $d$. The coupling phase $\varphi$ can be tuned to a nonzero value, breaking parity symmetry. When the propagation phase at the driving frequency matches $\varphi$, the giant system interacts only with the right-moving driving field (indicated by the right arrow), which is either reflected or transmitted, while it remains transparent to the left-moving input (indicated by the red arrow). In the former case, the transmitted light reveals distinctive characteristics, such as nontrivial statistical behaviors and the occurrence of dissipative phase transitions (DPTs).
By tuning the phase $\phi$ to match the propagation phase at a driving frequency $k_{i}$, corresponding to $\Omega_{\mathrm{L}}\left(-k_{i}\right)$ and $\Omega_{\mathrm{R}}\left(k_{i}\right)$, the giant system can be selectively decoupled from either the left- or right-moving driving field, effectively realizing the chiral light-emitter interaction. Additionally, reversing the direction of the input photons is equivalent to changing $\phi$ to $\phi$. Without loss of generality, in the following, we focus on the case with
$\phi=\left(2n+1\right)\mathrm{\pi}-\left(k_{0}+k_{i}\right)d,$ (3)
which results in $\Omega_{L}\left(-k_{i}\right)=0$, where $n$ is an integer, and study the chiral dynamics for right-moving inputs with frequency $k_{i}$.
II.1 Weak-driving limit
In the weak-driving limit, the dynamics are governed by the lowest-order behavior of $\Omega_{0}$ and can be linked to few-photon scattering processes Shi et al. (2015); Chang et al. (2016). For instance, reflection and transmission probabilities can be acquired by calculating the single-photon scattering processes, while second-order coherence can be inferred from the two-photon wavefunction. In the case of single-photon scattering, the reflection is reciprocal due to the $\mathcal{PT}$ symmetry Sakurai and Napolitano (2020); Yang and Chang (2024) of the system. Namely, the reflection coefficients satisfy $\left\langle 0\right|a_{\mathrm{R}k}Sa_{\mathrm{L}-k_{i}}^{\dagger}\left|0\right\rangle=\left\langle 0\right|a_{\mathrm{L}-k}Sa_{\mathrm{R}k_{i}}^{\dagger}\left|0\right\rangle$, where $|0\rangle$ is the vacuum state and the $S$ matrix is defined as $S=\lim_{T\rightarrow+\infty}^{\Omega_{0}\rightarrow 0}e^{iH_{0}T}e^{-i2HT}e^{iH_{0}T}$ Sakurai and Napolitano (2020). Consequently, the transmission probabilities for single left-moving and right-moving input photons are identical, since no other dissipation channels are present. Following the approach of Shi et al. (2015), where the bath field is integrated out using a path-integral method with the coherent-state representation (see Supplementary Note 1), we obtain the self-energy Fetter and Walecka (2012); Shi et al. (2016)
$\Sigma\left(k\right)$ $\equiv$ $\sum_{\lambda=\mathrm{L,R}}\int\frac{|V_{\lambda k}|^{2}\mathrm{d}k}{\omega-\varepsilon_{\lambda k}+i0^{+}}$ (4)
$=$ $-i2\gamma\left(1+\cos\phi e^{i\theta_{\mathrm{R}k}}\right),$
where $\gamma=2\mathrm{\pi}V_{0}^{2}$ is the decay rate accounting for cavity-bath coupling at a single point. The real part of $\Sigma\left(k\right)$ is the energy correction to the giant system, while its imaginary part gives to the decay rate. Consequently, the reflection and transmission coefficients with input and output momenta $k_{i}$ and $k_{\mathrm{f}}$, can be written as (see Supplementary Note 2) $\left\langle 0\right|a_{\mathrm{L}k_{\mathrm{f}}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle=r\left(k_{i}\right)\delta\left(k_{i}+k_{\mathrm{f}}\right)$ and $\left\langle 0\right|a_{Rk_{\mathrm{f}}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle=t\left(k_{i}\right)\delta\left(k_{i}-k_{\mathrm{f}}\right)$, where
$r\left(k\right)=-i\gamma\eta_{\mathrm{L}-k}^{\ast}\left(\phi\right)G\left(k\right)\eta_{\mathrm{R}k}\left(\phi\right)$ (5)
and
$t\left(k\right)=1-i\gamma\left|\eta_{\mathrm{R}k}\left(\phi\right)\right|^{2}G\left(k\right),$ (6)
with the Green function Fetter and Walecka (2012); Shi et al. (2015, 2016) $G\left(k\right)=\left[k-\Sigma\left(k\right)\right]^{-1}$.
Under the condition given by Eq. (3), the $\mathcal{PT}$ symmetry of the system ensures that the transmission probability for single right-moving photons is unity, indicating deterministic photon generation in the transmitted direction for both input orientations. However, the statistical properties of the right-moving transmitted photons can markedly differ from the trivial one for those ones moving to the left. Specifically, for two input photons with identical momentum $k_{i}$, the $S$-matrix element $S_{p_{2}p_{1},k_{i}k_{i}}\equiv\left\langle 0\right|a_{\mathrm{R}p_{2}}a_{\mathrm{R}p_{1}}Sa_{\mathrm{R}k_{i}}^{\dagger}a_{\mathrm{R}k_{i}}^{\dagger}\left|0\right\rangle$ governing the transmission of two photons with momenta $p_{1}$ and $p_{2}$ is (see Supplementary Note 3)
$S_{p_{2}p_{1},k_{i}k_{i}}$ $=$ $2t\left(k_{i}\right)^{2}\delta\left(k_{i}-p_{1}\right)\delta\left(k_{i}-p_{2}\right)$ (7)
$-i\Gamma_{p_{2}p_{1},k_{i}k_{i}}\left(\phi\right)\delta\left(2k_{i}-p_{1}-p_{2}\right)$
where
$\Gamma_{p_{2}p_{1},k_{i}k_{i}}=\frac{\gamma^{2}}{\pi}\eta_{\mathrm{R}p_{2}}^{\ast}\eta_{\mathrm{R}p_{1}}^{\ast}G\left(p_{1}\right)G\left(p_{2}\right)T_{S}G^{2}\left(k_{i}\right)\eta_{\mathrm{R}k_{i}}^{2}$ (8)
and
$T_{S}=\frac{U}{1-\frac{U}{k_{i}+i2\gamma}\sum_{mn}F_{mn}\left(k_{i}\right)\Theta\left(m-n\right)}$ (9)
depends on the input momentum $k_{i}$. The function
$F_{mn}\left(k\right)$ $=$ $\frac{e^{i\left(m-n\right)\left(k+i2\gamma\right)d}}{m!}\times$ (10)
$\left(\frac{-i\gamma\cos\phi e^{i\theta_{\mathrm{R}k}}}{k+i2\gamma}\right)^{m+n}\sum_{l=0}^{n}C_{mnl}\left(k\right),$
with $C_{mnl}\left(k\right)=\frac{\left(m+n-l\right)!}{\left(n-l\right)!l!}\left[-i2\left(m-n\right)\left(k+i2\gamma\right)d\right]^{l}$. In Eq. (7), the first term is resulted from the independent scattering of the two photons, while the second term represents the background fluorescence Shen and Fan (2007); Shi and Sun (2009).
The second-order correlation function
$g_{\lambda}^{\left(2\right)}\left(\tau\right)\equiv\frac{\left\langle a_{out}^{\lambda\dagger}\left(0\right)a_{out}^{\lambda\dagger}\left(\tau\right)a_{out}^{\lambda}\left(\tau\right)a_{out}^{\lambda}\left(0\right)\right\rangle}{\left\langle a_{out}^{\lambda\dagger}\left(0\right)a_{out}^{\lambda}\left(0\right)\right\rangle^{2}},$ (11)
where the output fields $a_{out}^{\lambda}\left(t\right)=\frac{1}{\sqrt{2\pi}}\int a_{\lambda k}\left(T\right)e^{-i\sigma_{\lambda}k\left(t-T\right)}\mathrm{d}k$, is connected to the two-photon wavefunction $w_{\lambda}\left(X_{1},X_{2}\right)$ with two photons at positions $X_{1}$ and $X_{2}$ as $g_{\lambda}^{\left(2\right)}\left(\tau\right)=\left|\frac{w_{\lambda}\left(\tau,0\right)}{w_{\lambda}\left(\tau\rightarrow\infty,0\right)}\right|^{2}$, while the wavefunction $w_{t}\left(X_{1},X_{2}\right)$ of two transmitted photons can be acquired by performing the Fourier transform of $S_{p_{2}p_{1},k_{2}k_{1}}$ as
$w_{t}\left(X_{1},X_{2}\right)=\frac{1}{2\pi}\int dp_{1}dp_{2}e^{ip_{1}X_{1}}e^{ip_{2}X_{2}}S_{p_{2}p_{1},k_{2}k_{1}}.$ (12)
Figure 2: Properties of the output photons in the weak-driving limit. (a) Second-order correlation function $g_{R}^{(2)}(0)$ of the transmitted light with $\phi=\pi/15$ and $U=\gamma$. (b) Same as (a), but for weak nonlinearity $U=0.1\gamma$. (c) $g_{R}^{(2)}(0)$ as a function of $k_{i}$ and $\phi$ in the strong dissipation limit with $U=0.01\gamma$ and $\gamma d=100.$ (d) Background fluorescence $\left|\Gamma_{p_{2}p_{1},k_{i}k_{i}}\right|$ with parameters from the rectangle in (c) shown in blue solid line. Here, in $\Gamma_{p_{2}p_{1},k_{i}k_{i}}$, $p_{2}=2k_{i}-p_{1}$. For comparison, $\left|\Gamma_{p_{2}p_{1},k_{i}k_{i}}\right|$ in the Markovian regime with $\gamma d=0.01$ is shown as the red dashed curve.
By fixing $\theta_{\mathrm{R}k_{i}}=\mathrm{\pi}-\phi$, we investigate the statistics of transmitted photons for right-moving inputs, varying both $k_{i}$ and $\phi$. The results are depicted in Fig. 2. When $U$ is comparable to the decay rate $\gamma$, the output photons can exhibit sub-Poissonian ($g_{\mathrm{R}}^{(2)}(0)<1$) or super-Poissonian ($g_{\mathrm{R}}^{(2)}(0)>1$) statistics for input momenta $k_{i}$ within a width approximately $\gamma$. In contrast to the single-point coupling scenario Shi et al. (2011), nontrivial statistical behaviors of transmitted photons persist even in the strong dissipation regime where $\gamma\gg U$, as shown in Fig. 2(b).
In the Makovian regime with $\gamma d\ll 1$, the second-order correlation function
$g_{\mathrm{R}}^{\left(2\right)}\left(0\right)=\left|\tilde{t}\left(k_{i}\right)^{2}+4\gamma_{0}^{2}\tilde{G}\left(k_{i}\right)^{2}-\frac{\gamma_{0}^{2}\tilde{G}\left(k_{i}\right)}{k_{i}-h}\right|^{2},$ (13)
[…]
Further tuning of $k_{i}$ and $\varphi$ reveals that the second-order correlation function $g_{\mathrm{R}}^{(2)}(0)$ varies in a wide range from $10^{-2}$ to $10^{4}$ (Fig. 2(c)). This nontrivial statistical behavior arises from photon interference between the two coupling points, resulting in an effective decay rate $\gamma_{k}\equiv$-Im[$\left[\Sigma(k)\right]=2\gamma\left(1+\cos\phi\cos\theta_{\mathrm{R}k}\right)$ for the $k$-th mode, which can be significantly smaller than $U$. These modes can be efficiently excited in the two-photon scattering process by matching its momentum $k$ with the real part of the self-energy, $\text{Re}\left[\Sigma(k)\right]=2\gamma\cos\phi\sin\theta_{\mathrm{R}k}$ or $2k_{i}$-Re[$\left[\Sigma(2k_{i}-k)\right]$. Fig. 2(d) illustrates this phenomenon with the background fluorescence Shen and Fan (2007); Shi and Sun (2009) $\Gamma_{p_{2}p_{1},k_{i}k_{i}}$ at the point indicated by the rectangle in Fig. 2(c), where $k_{i}=0.09\gamma$, $\phi=0.015\pi$, and $g_{\mathrm{R}}^{(2)}(0)=0.017$, corresponding to $\Sigma(k_{i})=(0.094-i0.0044)\gamma$. The blue solid curve in Fig. 2(d) shows highly populated modes centered around $0.094+2n\pi/d$ with a narrow width of $\gamma_{k}/\left(\gamma L\right)$, underscoring the nontrivial statistical behaviors of transmitted photons. In contrast, when $d$ decreases, the number of modes satisfying k=k=Re[$\left[\Sigma(k)\right]$ decrease, leading to less resonant peaks in the background fluorescence, together with larger width $\gamma_{k}$. Specifically, as shown in the red dashed curve, in the Makovian regime with $\gamma d=0.01$, the equation k=k=Re[$\left[\Sigma(k)\right]$ has only one resonant peak, resulting in different statistics of output photons.
We note that throughout the manuscript, the frequency of the cavity is set equal to the central frequency $k_{0}$ of the waveguide. In a gernal case, a detuning $\Delta_{c}$ between the cavity frequency and $k_{0}$ can be introduced. With this detuning, all occurrences of $k$ in the expressions—except in the phase factor $\theta_{\lambda k}$—should be replaced by $k-\Delta_{\mathrm{c}}$. By redefining $\tilde{k}=k-\Delta_{\mathrm{c}}$, the phase $\theta_{\mathrm{R}k}=\left(k_{0}+\tilde{k}_{i}+\Delta_{c}\right)d$, and the chirality condition in Eq. (3) is modified to
$\phi=\left(2n+1\right)\mathrm{\pi}-\left(k_{0}+\tilde{k}_{i}+\Delta_{c}\right)d,$ (14)
Since $\theta_{\mathrm{R}k}=\left(2n+1\right)\pi-\phi$, this redefinition allows all other expressions to retain their original form with $k$ replaced directly by $\tilde{k}$. Thus, when specifying $\tilde{k}_{i}$ as the input momentum, the results remain unchanged. For given values of $\phi$, $k_{0}$, and $d$, the chirality condition can be satisfied for different $\tilde{k}_{i}$ by tuning $\Delta_{\mathrm{c}}$.
II.2 Markovianity at $\phi=\pm\mathrm{\pi}/2$
When $\gamma d\ll 1$, the two-photon dynamics simplify to be Markovian (dashed-dotted line in Figs. 2(a) and (b)), and the system’s evolution can be approximatly described by a master equation and the input-output formalism Gardiner and Zoller (2004); Walls and Milburn (2008). Generally, this Markovian approximation breaks down for large distances $d$, as evidenced by the deviation in the second-order correlator $g_{\mathrm{R}}^{(2)}(0)$ between $\gamma d=0.01$ and $\gamma d=1,100$ shown in Figs. 2(a) and (b) (for fixed $\theta_{\mathrm{R}k_{i}}$, Markovian dynamics are independent of $d$). However, when $\phi=\pm\mathrm{\pi}/2$, the self-energy $\Sigma(k)$ becomes independent of $k$, and the evolution of the giant cavity’s density matrix $\rho$ can be exactly described by the master equation Shi et al. (2015) regardless of the distance $d$ and photon number:
$\partial_{t}\rho=-i\left[H_{S},\rho\right]+2\gamma\left(2b\rho b^{\dagger}-\left\{b^{\dagger}b,\rho\right\}\right),$ (15)
where $\rho$ is the density matrix of the cavity and $H_{S}=\frac{U}{2}b^{{\dagger}2}b^{2}+H_{\lambda D}$. In this scenario, the decay of the cavity due to the interference between right- and left-moving photons propagating between the two coupling points is completely canceled out.
This cancellation can also be understood by defining two modes
$\left(\begin{array}[]{c}A_{+k}\\ A_{-k}\end{array}\right)=U_{k}\left(\begin{array}[]{c}a_{L-k}\\ a_{Rk}\end{array}\right),$ (16)
where the unitary transformation $U_{k}=\frac{1}{\tilde{V}_{k}}\left(\begin{array}[]{cc}V_{\mathrm{L}-k}&V_{\mathrm{R}k}\\ V_{\mathrm{R}k}^{\ast}&-V_{\mathrm{L}-k}^{\ast}\end{array}\right)$, with $\tilde{V}_{k}=2V_{0}\sqrt{1+\cos\phi\cos\left(k_{0}+k\right)L}$. These two modes are degenerated and decoupled, and the cavity field only couples to $A_{+k}$ mode with the interaction in Eq. (2) rewritten as
$H_{1}=\int_{-\infty}^{+\infty}\tilde{V}_{k}\left(b^{{\dagger}}A_{+k}+\mathrm{H.c.}\right)\mathrm{d}k\mathrm{.}$ (17)
When $\phi=\pm\mathrm{\pi}/2$, $\tilde{V}_{k}$ is independent of $k$, leading to a constant decay rate $2\gamma$ contributed from the incoherent superposition of cavity photons decaying via the two coupling points. Note that this $k$-independence in $\tilde{V}_{k}$ with $\phi=\pm\pi/2$ is general and does not depend on the specific spectrum $\varepsilon_{\lambda k}$. However, the Markovian description is rigorously valid only for a linear spectrum. Furthermore, the properties of photons propagating in the waveguide can be acquired with the time-nonlocal input-output relation
$a_{\mathrm{out}}^{\lambda}\left(t\right)=a_{\mathrm{in}}^{\lambda}\left(t\right)-i\sqrt{\gamma}\left[e^{-i\phi}b\left(t\right)+e^{-i\theta_{\lambda 0}}b\left(t+\sigma_{\lambda}d\right)\right],$ (18)
where the input field $a_{\mathrm{in}}^{\lambda}\left(t\right)=\frac{1}{\sqrt{2\pi}}\int a_{\lambda k}\left(0\right)e^{-i\sigma_{\lambda}kt}\mathrm{d}k$. Note that unlike the giant system itself, the output field $a_{\mathrm{out}}^{\lambda}\left(t\right)$ is determined by the cavity field at two distinct times $t$ and $t\pm d$ even when $\phi=\pm\mathrm{\pi}/2$. This indicates that the non-Markovian characteristic is preserved in the properties of the output photons.
II.3 Non-Makovian dissipative phase transition
The master equation (15), along with the input-output formalism (18), enables a comprehensive exploration of the giant system’s dynamics and the characteristics of the output photons under arbitrarily strong driving fields. When $\phi=(k_{0}+k_{i})d=\mathrm{\pi}/2$, the giant system interacts exclusively with the right-moving driving field, leading to the emergence of a DPT in the strong-driving regime Drummond and Walls (1980); Drummond and Gardiner (1980); Bartolo et al. (2016). As depicted in Fig. 3(a), increasing the pump strength $\Omega_{0}$ results in a sudden increase in the cavity photon number $\langle b^{{\dagger}}b\rangle_{\mathrm{S}}$ around $\Omega_{0}\approx 5\gamma$, corresponding to a peak in the second-order correlator $g^{(2)}(0)\equiv\frac{\langle b^{{\dagger}2}b^{2}\rangle_{\mathrm{S}}}{\langle b^{{\dagger}}b\rangle_{\mathrm{S}}^{2}}$, where $\langle...\rangle_{\mathrm{S}}$ denotes the steady-state average. Here, we consider the Kerr nonlinearity $U=\gamma$ and incident momentum $k_{i}=10\gamma$. Using methods such as the complex P-representation formalism Walls and Milburn (2008); Drummond and Walls (1980); Drummond and Gardiner (1980); Bartolo et al. (2016) or the quantum absorber approach Stannigel et al. (2012); Roberts and Clerk (2020), one can analytically derive the steady-state solution of the master equation (15), revealing a first-order DPT in the thermodynamic limit where the Liouvillian gap (the real part of the second-largest eigenvalue of the Liouvillian) vanishes at the critical point Minganti et al. (2018).
This DPT-induced sudden change manifests not only in the cavity field but also in the reflected field, highlighting the non-Markovian effects induced by the giant system. At $T\rightarrow\infty$, the density $\rho_{L}\equiv\frac{1}{2\mathrm{\pi}}\int\langle a_{Lk_{2}}^{{\dagger}}(T)a_{Lk_{1}}(T)\rangle e^{i(k_{1}-k_{2})X}\mathrm{d}k_{1}\mathrm{d}k_{2}$ of left-moving output photons at position $X<0$ can be computed using the input-output relation (18) as
$\rho_{L}$ $=$ $\left\langle a_{\mathrm{out}}^{\mathrm{L}{\dagger}}\left(T+x\right)a_{\mathrm{out}}^{\mathrm{L}}\left(T+x\right)\right\rangle$ (19)
$=$ $2\gamma\left(\left\langle b^{{\dagger}}b\right\rangle_{\mathrm{S}}-\text{Im}\left[e^{ik_{0}d}\left\langle b^{{\dagger}}\left(d\right)b\left(0\right)\right\rangle_{\mathrm{S}}\right]\right).$
The DPT observed in the reflected field can differ significantly from that of the cavity field. For instance, under the conditions $\phi=\pm\pi/2$ and Eq. (3), the phase transition in the cavity field is independent of $d$. In contrast, the critical behavior in the reflected field is highly sensitive to $d$ due to the nonlocal correlation $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}$. When $d$ approaches zero, no reflected photons are observed, and thus no phase transition occurs in the reflected field. As $d$ increases, the DPT in the reflected field emerges, revealing the non-Markovian nature of the critical behavior in the output field.
Figure 3: Dissipative phase transition occurs only when the giant system is driven by right-moving input. (a) Photon number (blue solid line, left axis) in the cavity and the second-order correlation function (red line with circles, right axis), as functions of the driving strength (b) Photon number density (blue lines, left axis) of the reflected field and the two-point correlation (red lines with circles, right axis) for different distance $\gamma d=0.63$ (solid lines), $\gamma d=63$ (dashed lines), and $\gamma d=6300$ (dotted lines). Other parameters are $U=\gamma$ and $k_{i}=10\gamma$.
To quantitatively assess the impact of non-Markovian effect on the DPT, we set $k_{0}d=\pi/2$ and $d=2n\pi/k_{i}$ to satisfy the chirality condition in Eq. (3), and examine the DPT via $\rho_{L}$ for various values of $d$. As $d$ increases, the DPT in the reflected field becomes more pronounced, as illustrated in Fig. 3(b). Additionally, a peak emerges near the critical point where the Liouvillian gap reaches a minimum (vanishing only in the limit of infinite photon number), resulting in maximal correlations. $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}-\left|\left\langle b\right\rangle_{\mathrm{S}}\right|^{2}$. As $d$ further increases, the peak in $\rho_{L}$ approaches the critical point, and the magnitude of the correlator $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}-\left|\left\langle b\right\rangle_{\mathrm{S}}\right|^{2}$ diminishes because only the eigenstate with the smallest gap contributes significantly. Specifically, when $d$ exceeds the Liouvillian gap significantly, such as $d=2000\pi/k_{i}$, represented by the dotted lines in Fig. 3(b), no correlations are present and $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}$ approaches $\left|\left\langle b\right\rangle_{\mathrm{S}}\right|^{2}$. In this scenario, the photon density $\rho_{L}$ reaches its maximum since $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}$ deviates most significantly from $\left\langle b^{{\dagger}}b\right\rangle_{\mathrm{S}}$.
It is also worth noting that the reflected field behaves differently from the cavity field beyond the critical point. While the cavity photon number continues to grow with the driving amplitude $\Omega_{0}$, the reflected photon density $\rho_{L}$decreases after the critical point. This reduction arises from the suppression of the first-order correlator $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}-\left|\left\langle b\right\rangle_{\mathrm{S}}\right|^{2}$ after the transition. The diminishing correlations are further evidenced by the second-order correlation $g^{\left(2\right)}\left(0\right)$ approaching unity, as shown in Fig. 3(a). After the critical point, $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}\approx\left|\left\langle b\right\rangle_{S}\right|^{2}$ an $\rho_{L}\approx 2\gamma\left(\left\langle b^{{\dagger}}b\right\rangle_{\mathrm{S}}-\left|\left\langle b\right\rangle_{\mathrm{S}}\right|^{2}\right)$. Since both $\left\langle b^{{\dagger}}b\right\rangle_{\mathrm{S}}$ and $\left|\left\langle b\right\rangle_{\mathrm{S}}\right|^{2}$ increase with the driving amplitude $\Omega_{0}$, and the cavity field approaches a coherent state in the large $\Omega_{0}$ limit, the reflected photon density $\rho_{L}$ consequently decreases and eventually vanishes. These findings demonstrate that this system offers a valuable platform for exploring non-Markovian effects in DPTs.
II.4 Connection to experimental realizations
In this section, we present possible experimental implementations and discuss the impact of small deviations from the condition in Eq. (3) on the direction-dependent photon correlations and the nonreciprocal features of the dissipative phase transition.
II.4.1 Implementation
Our theoretical proposal can be implemented in superconducting circuit platforms. As demonstrated in a recent experiment Joshi et al. (2023), a giant system is coupled to a waveguide with a controllable phase difference between two spatially separated coupling points. As shown in Fig. 4(a), two sets of auxiliary resonators $C$ and $R$ are used. $C$ are tunable couplers, whose frequencies are modulated periodically via flux bias lines. $R$ are used to filter undesire emission into the waveguide. The phase difference $\phi_{\mathrm{L}}-\phi_{\mathrm{R}}$ corresponds to the phase $\phi$ in our model. The giant Kerr cavity can be realized as an LC circuit incorporating a nonlinear element like a Josephson junction Blais et al. (2021), realizing a finite Kerr nonlinearity $U$.
Figure 4: Schematics of the implementations of the theoretical model using superconducting circuits. In (a), four auxiliary resonators are employed. The frequencies of two of them—the couplers (CC)—are periodically modulated with a phase difference $\phi=\phi_{L}-\phi_{R}$. In (b), both the cavity frequency and its coupling to the waveguide at one end are periodically modulated. The coupling modulation is realized via a variable coupler composed of inductors and a SQUID.
[…]
$\tilde{H}_{S}=\left(\Delta+A\cos\Delta t\right)b^{{\dagger}}b+Ub^{{\dagger}2}b^{2}$ (20)
is the giant system’s Hamiltoian with modulation amplitude $A$ and
$\tilde{H}_{1}$ $=$ $\int_{-\infty}^{+\infty}\mathrm{d}k[V_{1}\left(a_{\mathrm{L}k}+a_{\mathrm{R}k}\right)b^{{\dagger}}\cos\left(\Delta t+\phi^{\prime}\right)+$ (21)
$V_{2}\left(a_{\mathrm{L}k}e^{-i\left(k_{0}-k\right)L}+a_{\mathrm{R}k}e^{i\left(k_{0}+k\right)L}\right)b^{{\dagger}}+\mathrm{H.c.}$
is its linear coupling to the waveguide, with coupling constants $V_{1}$ and $V_{2}$. In the rotating frame with respect to $\tilde{H}_{0}\left(t\right)=\left(\Delta+A\cos\Delta t\right)b^{{\dagger}}b$, the Hamitltonian becomes
$H=U^{\dagger}\left(t,0\right)\tilde{H}\left(t\right)U\left(t,0\right)-\tilde{H}_{0}\left(t\right)\approx H_{0}+H_{1},$ (22)
where the time evolution operator $U\left(t,0\right)=\mathcal{T}e^{-i\int_{0}^{t}\tilde{H}_{0}\left(t^{\prime}\right)dt^{\prime}}=e^{-i\left(\Delta t+\frac{A}{\Delta}\sin\Delta t\right)b^{{\dagger}}b}$. Here, we have ignored the fast oscillation terms assuming that $\left|V_{1,2}J_{n}\left(\frac{A}{\Delta}\right)\right|^{2}\ll\left|\Delta\right|$, and $V$ and $\phi$ are connected to the modulation parameters as
$V$ $=$ $\frac{V_{1}}{2}\left|J_{2}\left(\frac{A}{\Delta}\right)e^{i\phi^{\prime}}+J_{0}\left(\frac{A}{\Delta}\right)e^{-i\phi^{\prime}}\right|$ (23)
$=$ $V_{2}J_{1}\left(\frac{A}{\Delta}\right)$
and
$\phi=\text{Arg}\left(J_{2}\left(\frac{A}{\Delta}\right)e^{i\phi^{\prime}}+J_{0}\left(\frac{A}{\Delta}\right)e^{-i\phi^{\prime}}\right).$ (24)
II.4.2 Deviations from the chirality condition in (3)
When the condition in Eq. (3) is not fulfilled, the properties of the transmitted photons under right-moving driving are altered, and the behavior of transmitted photons under left-moving driving becomes nontrivial, since the giant system can now be driven by the left-moving input. For a small phase deviation $\Delta\phi$ from the value specified in Eq. (3), the second-order correlation functions $g_{R,L}^{\left(2\right)}\left(0\right)$ for the transmitted photons are shown in Figs. 5(a) and (b). The phase is set to $\phi=\mathrm{\pi}/15+\Delta\phi$, with other parameters the same as in Figs. 2(a) and (b).
Compare to the cases in Figs. 2(a) and (b) with right-moving driving, the transmission probability is slight less than $1$, and the second-order correlation function $g_{R}^{\left(2\right)}\left(0\right)$ shows only minor deviation for small $\Delta\phi=\pi/100$ and $\pi/50$. Meanwhile, $g_{L}^{\left(2\right)}\left(0\right)$ under left-moving driving exhibits nontrivial statistics, with $g_{L}^{\left(2\right)}\left(0\right)$ slightly deviating from 1. Notably, the deviations in $g_{R,L}^{\left(2\right)}\left(0\right)$ diminish for smaller $U$, because only modes with smaller effective decay rate $\gamma_{k}$ contribute to nontrivial second-order correlation functions. For smaller $U$, fewer of these modes are effectively excited, resulting in a reduced sensitivity to $\Delta\phi$. Therefore, for sufficiently small $\Delta\phi$, $g_{L}^{\left(2\right)}\left(0\right)$ for left-moving driving remains approximately 1, and nontrivial photon statistics are generated predominantly for a single incident direction.
Figure 5: Deviations from the chirality condition. Second-order correlation function $g_{R,L}^{(2)}(0)$ of the transmitted photons in the weak-driving limit, and the density of the reflected photons in the strong-driving limit, when the chirality condition is slightly violated. The driving field is right-moving in (a) and (c), and left-moving in (b) and (d). In (a) and (b), the distance is $d=100/\gamma$, with $U=\gamma$ (blue solid and red dashed curves) and $0.1\gamma$ (yellow dotted and purple dash-dotted curves). The phase deviations are $\Delta\phi=\pi/100$ (blue solid and yellow dotted curves) and $\pi/50$ (red dashed and purple dash-dotted curves). The inset in (b) shows the transmission probability, with the legend consistent with that in (a) and (b). In (c) and (d), the phase is fixed at $\phi=\pi/2$, the distance is $d=6300/\gamma$, and different deviations of the propagation phase are considered. Other parameters are the same as in Fig. 3.
We now consider the impact of a small deviation from Eq. (3) on the DPT under the strong-diving limit. To ensure Markovian dynamics in the giant system, $\phi$ is set to be $\mathrm{\pi}/2$, and a small deviation $\Delta\theta$ is introduced, making the propagation phase $\theta_{\mathrm{R}k_{i}}=\mathrm{\pi}/2+\Delta\theta$. With other parameters identical to those in Fig. 3, the density of reflected photons under right-moving driving shows negligible change, as illustrated in Fig. 5(c). Under left-moving driving, the reflected photon density $\rho_{R}$ also exhibits a DPT, but at a critical pumping amplitude $\Omega_{0}$ that is more than 10 times larger, as shown in Fig. 5(d). This is because the effective driving amplitude $\Omega_{L}\left(k_{i}\right)\ll\Omega_{R}\left(k_{i}\right)$. As a result, if we focus on the critical behavior within the driving amplitude range around the critical point for right-moving driving, the DPT remains nonreciprocal.
III Conclusion
In summary, we have explored the non-Markovian multiphoton dynamics of a driven giant system under chiral coupling conditions, achieved by matching the phase difference $\phi$ between coupling points to the propagation phase of the input field. Focusing on a nonlinear giant cavity coupled unidirectionally to a waveguide, we demonstrated that in the weak-driving (few-photon) regime, strongly nonclassical photon statistics can be deterministically generated—even in the presence of strong dissipation and weak nonlinearity. This arises from coherent interference between forward- and backward-propagating components within the nonlocal coupling region and is tunable via the coupling-point separation, reflecting the intrinsic non-Markovianity of the system. At the special point $\phi=\pm\mathrm{\pi}/2$, interference between coupling points vanishes, rendering the system’s internal dynamics exactly Markovian. Nonetheless, the output fields remain non-Markovian due to the extended coupling structure. In this regime, we uncover a nonreciprocal dissipative phase transition in the reflected field, governed by nonlocal photon correlations. The physical implementations are presented, along with a discussion of the impact of small deviations from the chirality condition. Our study can be extended to casese with $N$ coupling legs (see Supplementary Note 4). The findings offer a robust platform for engineering chiral, non-Markovian multiphoton states in driven-dissipative settings. This framework is broadly applicable to a variety of nonlocally coupled quantum systems, including giant cavities interfaced with atomic ensembles and central spin models coupled to spin chains, providing a foundation for scalable quantum state control in open quantum networks.
IV Methods
IV.1 Numerical simulations
The dissipative phase transitions shown in Figs. 3 and 5 are obtained by solving the master equation (15) and using the input-output formalism (18). the master equation (15) is solved in the superspace, where the density matrix $\rho$ is reshaped as a column vector $\vec{\rho}$ of dimension $D^{2}$. Here, $D$ is the Hilber space dimension, with D1D-1 corresponding to the cutoff of the boson number for mode $b$. In the superspace, the Liouvillian $\mathcal{L}$, defined as
$\mathcal{L}\rho\equiv-i\left[H_{S},\rho\right]+2\gamma\left(2b\rho b^{\dagger}-\left\{b^{\dagger}b,\rho\right\}\right),$ (25)
is represented as a $D^{2}\times D^{2}$ matrix
$\mathcal{L}_{M}$ $=$ $-i\left(H_{S}\otimes I_{D}-I_{D}\otimes\left(H_{S}\right)^{T}\right)+$ (26)
$2\gamma\left(2b\otimes\left(b^{\dagger}\right)^{T}-b^{\dagger}b\otimes I_{D}-I_{D}\otimes b^{\dagger}b\right),$
where $I_{D}$ is the $D$-dimensional identity matrix, and $\left(\bullet\right)^{T}$ denotes the matrix transpose.
The steady state solution $\vec{\rho}_{ss}$ of the master equation can be obtained by diagonalizing the Liouvillian matrix $\mathcal{L}_{M}$ and identifying the zero-eigenvalue state. The steady state must satisfy the normalization condition $u_{0}\vec{\rho}_{ss}=1$, which follows from Tr[$\left[\rho\right]=1$, where $u_{0}$ is generated by reshaping the $D$-dimensional identity matrix $I_{D}$ into a $1\times D^{2}$ row vector. Alternatively, $\vec{\rho}_{ss}$ can be found from the equation $\mathcal{L}_{M}\vec{\rho}_{ss}=0$. Note that $\mathcal{L}_{M}$ is not reversible (its rank is less than $D^{2})$, and that is why it has zero eigenvalues. Physically, this reflects Tr[$\left[\rho\right]=1$, which implies Tr[$\left[\mathcal{L}\rho\right]=0$. To incorporate the normalization condition Tr[$\left[\rho\right]=1$, one replaces the row corresponding to the diagonal-element of $\vec{\rho}$ ($1,D+1,2D+2,...,D^{2}$)with $u_{0}$. Without loss of generality, we replace the last row of $\mathcal{L}_{M}$ with $u_{0},$ forming a modified matrix $\mathcal{L}_{M}^{\prime}$. The steady state is then given by $\vec{\rho}_{ss}=\left(\mathcal{L}_{M}^{\prime}\right)^{-1}v_{0}$, where $v_{0}$ is a $D^{2}\times 1$ column vector with $1$ in the last element and zeros elsewhere.
Once the steady state $\vec{\rho}_{ss}$ is obtained, the long-term-average of any system operator can be calculated. For example, the mean value photon number value is given by
$\langle b^{{\dagger}}b\rangle_{S}=u_{0}\left(b^{{\dagger}}b\otimes I_{D}\right)\vec{\rho}_{ss}.$ (27)
Multi-time correlations function can be evaluated using the quantum regression theorem Gardiner and Zoller (2004); Walls and Milburn (2008). For instance, the two-time correlation function in Eq. (19) can be computed as
$\left\langle b^{{\dagger}}\left(d\right)b\left(0\right)\right\rangle=u_{0}\left(b^{{\dagger}}\otimes I_{D}\right)e^{\mathcal{L}_{M}d}\left(b\otimes I_{D}\right)\vec{\rho}_{ss}.$ (28)
V Data availability
Data that support the findings of this study are available from the corresponding author upon reasonable request.
VI Code availability
The codes used in this study are available from the corresponding author upon reasonable request.
VII Acknowledgements
Acknowledgements.
The author would like to thank Carlos Navarrete-Benlloch for fruitful discussions on the physical implementation.
Supplemental information for “Non-Markovian Multiphoton Chiral Dynamics with Giant Systems ”
Supplementary Note1: S-matrix in coherent-state presentation
We first rewrite the $S$ matrix for reflection and transmission photons in forms of the two modes $A_{\pm k}$ defined in Eq. (11) as
$\left(\begin{array}[]{c}A_{+k}\\ A_{-k}\end{array}\right)=U_{k}\left(\begin{array}[]{c}a_{L-k}\\ a_{Rk}\end{array}\right).$ (S1)
Since only the modes $A_{+k}$ couple to the giant system, the transmission amplitude $\left\langle 0\right|a_{Rk_{f}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle$ for a right-moving photon with input momentum $k_{i}$ and an output photon with momentum $k_{f}$ is
$\left\langle 0\right|a_{Rk_{f}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle$ $=$ $\left|U_{k_{i},22}\right|^{2}\delta\left(k_{i}-k_{f}\right)+\left|U_{k_{i},12}\right|^{2}\left\langle 0\right|A_{k_{f}+}SA_{k_{i}+}^{\dagger}\left|0\right\rangle$ (S2)
$=$ $\left|U_{k_{i},22}\right|^{2}\delta\left(k_{i}-k_{f}\right)+\left|U_{k_{i},12}\right|^{2}\left\langle 0\right|A_{k_{f}+}e^{-i2HT}A_{k_{i}+}^{\dagger}\left|0\right\rangle e^{i\left(k_{i}+k_{f}\right)T},$
where $T\rightarrow+\infty$. Similarly, the reflection amplitude
$\left\langle 0\right|a_{Lk_{f}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle=U_{-k_{f},21}^{\ast}U_{k_{i},22}\delta\left(k_{i}+k_{f}\right)+U_{-k_{f},11}^{\ast}U_{k_{i},12}\left\langle 0\right|A_{-k_{f}+}e^{-2HT}A_{k_{i}+}^{\dagger}\left|0\right\rangle e^{ik_{i}T}e^{-ik_{f}T}.$ (S3)
For two photons, the amplitude for right-moving incident photons with momenta $k_{1}$ and $k_{2}$, and transmitted photons with momenta $p_{1}$ and $p_{2}$, is
$\left\langle 0\right|a_{Rp_{1}}a_{Rp_{2}}Sa_{Rk_{2}}^{\dagger}a_{Rk_{1}}^{\dagger}\left|0\right\rangle$ (S4)
$=$ $U_{p_{2},22}^{\ast}U_{p_{1},22}^{\ast}U_{k_{2},22}U_{k_{1},22}\left[\delta\left(k_{1}-p_{1}\right)\delta\left(k_{2}-p_{2}\right)+\left(k_{1}\longleftrightarrow k_{2}\right)\right]$
$+\{U_{p_{2},22}^{\ast}U_{p_{1},12}^{\ast}U_{k_{2},22}U_{k_{1},12}\left\langle 0\right|A_{p_{1}+}e^{-i2HT}A_{k_{1}+}^{\dagger}\left|0\right\rangle e^{i\left(k_{1}+p_{1}\right)T}\delta\left(k_{2}-p_{2}\right)$
$+\left(k_{1}\longleftrightarrow k_{2}\right)+\left(p_{1}\longleftrightarrow p_{2}\right)+\left(k_{1}\longleftrightarrow k_{2},p_{1}\longleftrightarrow p_{2}\right)\}$
$+U_{p_{2},12}^{\ast}U_{p_{1},12}^{\ast}U_{k_{2},12}U_{k_{1},12}\left\langle 0\right|A_{p_{2}+}A_{p_{1}+}e^{-i2HT}A_{k_{2}+}^{\dagger}A_{k_{1}+}^{\dagger}\left|0\right\rangle e^{i\left(k_{1}+k_{2}+p_{1}+p_{2}\right)T}.$
Therefore, instead of the original left- and right-moving modes $a_{Lk}$ and $a_{L-k}$, we can obtain the S-matrix by considering only the $A_{k+}$ mode while the $A_{k-}$ mode is decoupled from the giant system. Here, we have assumed that in the initial and final states, the giant system is in vacuum, which is the case we focus on in the weak-driving limit.
To calculate the S-matrix ultilizing the path-integral method, we write it in the basis of coherent-states. The S-matrix for $N$ input photons with momenta $k_{i_{1}}k_{i_{2}},...,k_{i_{N}}$ and $N$ output photons with momenta $k_{f_{1}}k_{f_{2}},...,k_{f_{N}}$ envolves the calculation of
$\left\langle k_{f_{1}}k_{f_{2}}...k_{f_{N}}\right|e^{-i2HT}\left|k_{i_{1}}k_{i_{2}}...k_{i_{N}}\right\rangle,$ (S5)
where $\left|k_{1}k_{2}...k_{N}\right\rangle=A_{k_{N}+}^{\dagger}...A_{k_{2}+}^{\dagger}A_{k_{1}+}^{\dagger}\left|0\right\rangle$. Introducing an unnormalized coherent state Walls and Milburn (2008)
$\left|\alpha_{k}\right\rangle=\sum_{n_{k}=0}\frac{\alpha_{k}^{n}}{\sqrt{n_{k}!}}\left|n_{k}\right\rangle,$ (S6)
for the $k$-th mode $A_{k+}^{\dagger}$, the the evolution term $\left\langle k_{f_{1}}k_{f_{2}}...k_{f_{N}}\right|e^{-i2HT}\left|k_{i_{1}}k_{i_{2}}...k_{i_{N}}\right\rangle$ can be written as
$\left\langle k_{f_{1}}k_{f_{2}}...k_{f_{N}}\right|e^{-i2HT}\left|k_{i_{1}}k_{i_{2}}...k_{i_{N}}\right\rangle$ (S7)
$=$ $\lim_{\alpha_{k_{i_{1}}}\rightarrow 0}\partial_{\alpha_{k_{i_{1}}}}\lim_{\alpha_{k_{i_{2}}}\rightarrow 0}\partial_{\alpha_{k_{i_{2}}}}...\lim_{\alpha_{k_{i_{N}}}\rightarrow 0}\partial_{\alpha_{k_{i_{N}}}}\lim_{\alpha_{k_{f_{1}}^{\ast}}\rightarrow 0}\partial_{\alpha_{k_{f_{1}}^{\ast}}}\lim_{\alpha_{k_{f_{2}}^{\ast}}\rightarrow 0}\partial_{\alpha_{k_{f_{2}}^{\ast}}}...\lim_{\alpha_{k_{f_{N}}^{\ast}}\rightarrow 0}\partial_{\alpha_{k_{f_{N}}^{\ast}}}\mathcal{A},$
where $\mathcal{A}\equiv\left\langle\alpha_{k_{f_{1}}}\alpha_{k_{f_{2}}}...\alpha_{k_{f_{N}}}\right|e^{-i2HT}\left|\alpha_{k_{i_{1}}}\alpha_{k_{i_{2}}}...\alpha_{k_{i_{N}}}\right\rangle$. Eq. (S7) connects the S-matrix in the fock-state basis to the coherent-state representaion. Note that the overlap of two reduced coherent states is
$\left\langle\beta\right.\left|\alpha\right\rangle=e^{\alpha\beta^{\ast}}$ (S8)
and the completeness property of the reduced coherent states reads
$\int\mathrm{d}^{2}\alpha\text{ }\frac{e^{-\left|\alpha\right|^{2}}}{\pi}\left|\alpha\right\rangle\left\langle\alpha\right|=1.$ (S9)
The amplitude $\mathcal{A}$ in the path-integral form is Shi et al. (2015)
$\mathcal{A}=\int D\left[\beta,\beta^{\ast},\alpha_{k},\alpha_{k}^{\ast}\right]e^{\sum_{k}\alpha_{k}^{\ast}\left(-T\right)\alpha_{k}\left(-T\right)}\exp\left[i\int_{-T}^{T}\mathrm{d}t\text{ }\left[\mathcal{L}_{w}\left(t\right)+\mathcal{L}_{c}\left(t\right)\right]\right],$ (S10)
where
$D\left[\beta,\beta^{\ast},\alpha_{k},\alpha_{k}^{\ast}\right]=\prod\limits_{t}\left[\frac{d\beta\left(t\right)d\beta^{\ast}\left(t\right)}{\pi}\prod\limits_{k}\frac{d\alpha_{k}\left(t\right)d\alpha_{k}^{\ast}\left(t\right)}{\pi}\right],$ (S11)
is the measure
$\mathcal{L}_{w}\left(t\right)=-i\sum_{k}\dot{\alpha}_{k}^{\ast}\left(t\right)\alpha_{k}\left(t\right)-\sum_{k}\left[k\alpha_{k}^{\ast}\left(t\right)\alpha_{k}\left(t\right)+\tilde{V}_{k}\beta^{\ast}\left(t\right)\alpha_{k}\left(t\right)+\tilde{V}_{k}\alpha_{k}^{\ast}\left(t\right)\beta\left(t\right)\right],$ (S12)
is the Lagrangian for the waveguide and the emitter-waveguide interaction, and
$\mathcal{L}_{c}\left(t\right)=-i\dot{\beta}^{\ast}\left(t\right)\beta\left(t\right)-\frac{U}{2}\beta^{\ast}\left(t\right)\beta^{\ast}\left(t\right)\beta\left(t\right)\beta\left(t\right),$ (S13)
is the Lagrangian for the giant cavity. At the the initial and final moments, $\beta\left(-T\right)=\beta\left(T\right)=0$, $\left\{\alpha_{k}\left(-T\right)\right\}=\left\{\alpha_{k_{i}}\right\}$, and $\left\{\alpha_{k}\left(T\right)\right\}=\left\{\alpha_{k_{f}}\right\}$. Following the saddle-point method Shi et al. (2015), we first calculate the classical trajectory given by $\delta_{\alpha_{k}}\mathcal{L}_{ph}-\partial_{t}\frac{\delta\mathcal{L}_{ph}}{\delta_{\dot{\alpha}_{k}}}=0$ and obtain
$\alpha_{k,cl}^{\ast}\left(t\right)=\alpha_{k}^{\ast}\left(T\right)e^{-ik\left(T-t\right)}-iV_{k}\int_{t}^{T}\beta^{\ast}\left(t^{\prime}\right)e^{-ik\left(t^{\prime}-t\right)}\mathrm{d}t^{\prime}.$ (S14)
Consequently, the waveguide modes can be integrated out and the amplitude $\mathcal{A}$ becomes
$\mathcal{A}$ $=$ $\int D\left[\beta,\beta^{\ast}\right]\exp\left[i\int_{-T}^{T}\mathrm{d}t\text{ }\mathcal{L}_{eff}\left(t\right)\right]\times$ (S15)
$\exp\left[\alpha_{k_{f}}^{\ast}\alpha_{k_{i}}e^{-i2k_{i}T}\delta\left(k_{i}-k_{f}\right)-i\tilde{V}_{k_{i}}\alpha_{k_{i}}\int_{-T}^{T}\mathrm{d}t\text{ }\beta^{\ast}\left(t\right)e^{-ik_{i}\left(t+T\right)}-i\tilde{V}_{k_{f}}\alpha_{k_{f}}^{\ast}\int_{-T}^{T}\mathrm{d}t\text{ }e^{-ik_{f}\left(T-t\right)}\beta\left(t\right)\right].$
where
$\mathcal{L}_{eff}\left(t\right)=\mathcal{L}_{c}\left(t\right)+i\int_{t}^{T}\mathrm{d}t^{\prime}\int\mathrm{d}k\text{ }\tilde{V}_{k}^{2}e^{-ik\left(t^{\prime}-t\right)}\beta^{\ast}\left(t^{\prime}\right)\beta\left(t\right)$ (S16)
is the effective Lagrangian for the giant system. We will see in the following that the second term in Eq. (S16) will result in decay and energy correction to the giant system.
Supplementary Note 2: Single-photon scattering
The the S-matrix element $\left\langle 0\right|A_{k_{f}+}SA_{k_{i}+}^{\dagger}\left|0\right\rangle$ for an incident photon with momentum $k_{i}$ and outgoing one with momentum $k_{f}$ is
$\left\langle 0\right|A_{k_{f}+}SA_{k_{i}+}^{\dagger}\left|0\right\rangle$ (S17)
$=$ $e^{i\left(k_{i}+k_{f}\right)T}\lim_{\left\{\alpha_{k}\right\}\rightarrow 0}\partial_{\alpha_{k_{i}}}\partial_{\alpha_{k_{f}}^{\ast}}\mathcal{A}$
$=$ $\int D\left[\beta,\beta^{\ast}\right]\left[\delta\left(k_{i}-k_{f}\right)-\tilde{V}_{k_{i}}\tilde{V}_{k_{f}}\int_{-T}^{T}\mathrm{d}t\mathrm{d}t^{\prime}e^{-ik_{i}t^{\prime}}e^{ik_{f}t}\beta^{\ast}\left(t^{\prime}\right)\beta\left(t\right)\right]\exp\left[i\int_{-T}^{T}\mathrm{d}t\text{ }\mathcal{L}_{eff}\left(t\right)\right].$
Denoting the Fourier transform of $\beta\left(t\right)$ as
$\beta\left(t\right)=\frac{1}{\sqrt{2\pi}}\int\tilde{\beta}\left(\omega\right)e^{-i\omega t}\mathrm{d}\omega,$ (S18)
we acquire
$\left\langle 0\right|A_{k_{f}+}SA_{k_{i}+}^{\dagger}\left|0\right\rangle$ (S19)
$=$ $\delta\left(k_{i}-k_{f}\right)-i2\pi\tilde{V}_{k_{i}}\tilde{V}_{k_{f}}\int D\left[\tilde{\beta},\tilde{\beta}^{\ast}\right]\tilde{\beta}^{\ast}\left(k_{i}\right)\tilde{\beta}\left(k_{f}\right)\exp\left[\int\mathrm{d}\omega\left[\omega-\Sigma\left(\omega\right)\right]\tilde{\beta}^{\ast}\left(\omega\right)\tilde{\beta}\left(\omega\right)\right]$
$=$ $\delta\left(k_{i}-k_{f}\right)-i2\pi\tilde{V}_{k_{i}}^{2}G\left(k_{i}\right)\delta\left(k_{i}-k_{f}\right).$
where the Green’s function
$G\left(\omega\right)=\frac{1}{\omega-\Sigma\left(\omega\right)}$ (S20)
and the self energy
$\Sigma\left(\omega\right)\equiv\int\mathrm{d}k\frac{\tilde{V}_{k}^{2}}{\omega-k+i0^{+}}=-i2\gamma\left(1+\cos\phi e^{i\left(k_{0}+\omega\right)d}\right).$ (S21)
[…]
Consequently, the single-photon transmission amplitude is
$\left\langle 0\right|a_{Rk_{f}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle=\delta\left(k_{i}-k_{f}\right)-i2\gamma\left|e^{i\phi}+e^{i\left(k_{0}+k_{i}\right)d}\right|^{2}G\left(k_{i}\right)\delta\left(k_{i}-k_{f}\right)$ (S22)
and the reflection amplitude is
$\left\langle 0\right|a_{Lk_{f}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle=-i\gamma\left(e^{-i\phi}+e^{i\left(k_{0}+k_{i}\right)L}\right)\left(e^{i\phi}+e^{i\left(k_{0}+k_{i}\right)d}\right)G\left(k_{i}\right)\delta\left(k_{i}+k_{f}\right).$ (S23)
Supplementary Note 3: Two-photon scattering
For two incident photons with momenta $k_{1}$ and $k_{2}$, the S-matrix reads
$\left\langle 0\right|A_{p_{2}+}A_{p_{1}+}SA_{k_{2}+}^{\dagger}A_{k_{1}+}^{\dagger}\left|0\right\rangle$ (S24)
$=$ $e^{i\left(k_{1}+k_{2}+p_{1}+p_{2}\right)T}\lim_{\left\{\alpha_{k}\right\},\left\{\alpha_{p}\right\}\rightarrow 0}\partial_{\alpha_{p_{2}}^{\ast}}\partial_{\alpha_{p_{1}}^{\ast}}\partial_{\alpha_{k_{2}}}\partial_{\alpha_{k_{1}}}\mathcal{A}$
$=$ $\left[1-i2\pi\left(\tilde{V}_{k_{2}}^{2}G\left(k_{2}\right)+\tilde{V}_{k_{1}}^{2}G\left(k_{1}\right)\right)\right]\delta\left(k_{1}-p_{1}\right)\delta\left(k_{2}-p_{2}\right)+\left(k_{1}\leftrightarrow k_{2}\right)+B,$
which reveals two processes: the independent scattering of the two photons depicted by the first term in the last line of Eq. (S24), and the rest
$B$ $\equiv$ $\int D\left[\beta,\beta^{\ast}\right]\tilde{V}_{p_{2}}\tilde{V}_{p_{1}}\tilde{V}_{k_{2}}\tilde{V}_{k_{1}}\exp\left[i\int_{-T}^{T}\mathrm{d}t\text{ }\mathcal{L}_{eff}\left(t\right)\right]\times$ (S25)
$\int_{-T}^{T}\prod\limits_{j=1}^{4}\mathrm{d}t_{j}\text{ }\beta\left(t_{4}\right)e^{ip_{2}t_{4}}\beta\left(t_{3}\right)e^{ip_{1}t_{3}}\beta^{\ast}\left(t_{2}\right)e^{-ik_{2}t_{2}}\beta^{\ast}\left(t_{1}\right)e^{-ik_{1}t_{1}}.$
corresponds to the background fluorescence. Under the Fourier transform of $\beta\left(t\right)$ and the Dyson expansion Fetter and Walecka (2012), $B$ becomes
$B$ $=$ $-\tilde{V}_{p_{2}}\tilde{V}_{p_{1}}\tilde{V}_{k_{2}}\tilde{V}_{k_{1}}\left(2\pi\right)^{2}G\left(k_{1}\right)G\left(k_{2}\right)\left[\delta\left(k_{1}-p_{1}\right)\delta\left(k_{2}-p_{2}\right)+\delta\left(k_{1}-p_{2}\right)\delta\left(k_{2}-p_{1}\right)\right]$ (S26)
$-4i\pi\tilde{V}_{p_{2}}\tilde{V}_{p_{1}}\tilde{V}_{k_{2}}\tilde{V}_{k_{1}}G\left(k_{1}\right)^{-1}G\left(k_{2}\right)^{-1}T_{S}G\left(p_{1}\right)^{-1}G\left(p_{2}\right)^{-1}\delta\left(k_{1}+k_{2}-p_{1}-p_{2}\right),$
where the the giant system’s $T$-matrix can be exactly calculated as
$T_{S}=\frac{U}{1-i\frac{U}{2\pi}\int\mathrm{d}\omega G\left(\omega\right)G\left(k_{1}+k_{2}-\omega\right)}.$ (S27)
[…]
$\int\mathrm{d}\omega G\left(\omega\right)G\left(k_{1}+k_{2}-\omega\right)$ (S28)
$=$ $\int d\omega\sum_{mn}\frac{e^{im\left(k_{0}+\omega\right)L}\left(-i2\gamma\cos\phi\right)^{m}}{\left(\omega+i2\gamma\right)^{m+1}}\frac{e^{in\left(k_{0}+k_{1}+k_{2}-\omega\right)L}\left(-i2\gamma\cos\phi\right)^{n}}{\left(k_{1}+k_{2}-\omega+i2\gamma\right)^{n+1}}$
$=$ $\frac{2\pi}{i\bar{k}_{i}-2\gamma}\sum_{mn}F_{mn}\left(\bar{k}_{i}\right)\Theta\left(m-n\right).$
where $\bar{k}_{i}=\frac{k_{1}+k_{2}}{2}$ is the average momentum of the incident photons, and $F_{mn}\left(k\right)$ is defined as
$F_{mn}\left(k\right)=\frac{e^{\left(m-n\right)\left(ik-2\gamma\right)d}}{m!}\left(\frac{-i\gamma\cos\phi e^{i\theta_{Rk}}}{k+i2\gamma}\right)^{m+n}\sum_{l=0}^{n}C_{mnl}\left(k\right),$ (S29)
with
$C_{mnl}\left(k\right)=\frac{\left(m+n-l\right)!}{\left(n-l\right)!l!}\left[-i2\left(m-n\right)\left(k+i2\gamma\right)d\right]^{l}.$ (S30)
With Eq. (S4), we can obtain the S-matrix $S_{p_{2}p_{1},k_{i}k_{i}}$ for two transmitted photons shown in Eq. (7).
Performing the Fourier transform to $S_{p_{2}p_{1},k_{2}k_{1}}$, the wavefunction $w_{t}\left(X_{1},X_{2}\right)$ for two transmitted photons at positions $X_{1}$ and $X_{2}$ is
$w_{t}\left(X_{1},X_{2}\right)$ (S31)
$=$ $\frac{1}{2\pi}t\left(k_{1}\right)t\left(k_{2}\right)\left(e^{ik_{2}X_{1}}e^{ik_{1}X_{2}}+e^{ik_{1}X_{1}}e^{ik_{2}X_{2}}\right)$
$+i\frac{\gamma^{2}}{2\pi}\left[e^{i\phi}+e^{i\left(k_{0}+k_{1}\right)d}\right]\left[e^{i\phi}+e^{i\left(k_{0}+k_{2}\right)d}\right]TG\left(k_{1}\right)G\left(k_{2}\right)F_{t}\left(X_{1},X_{2}\right)$
where the correlated part
$F_{t}\left(X_{1},X_{2}\right)$ (S32)
$\equiv$ $\frac{1}{\pi}\int\mathrm{d}pe^{ipX_{1}}e^{i\left(k_{1}+k_{2}-p\right)X_{2}}\left(e^{-i\phi}+e^{-i\left(k_{0}+p\right)d}\right)\left(e^{-i\phi}+e^{-i\left(k_{0}+2\bar{k}_{i}-p\right)d}\right)G\left(p\right)G\left(k_{1}+k_{2}-p\right)$
$=$ $\frac{1}{\pi}e^{i2\bar{k}_{i}X_{2}}\sum_{mn}\left(-i2\gamma\cos\phi\right)^{m+n}\left(-\right)^{n+1}e^{imk_{0}L}e^{in\left(k_{0}+2\bar{k}_{i}\right)L}\int dp\frac{e^{ipX_{mn}}\left(e^{-i\phi}+e^{-i\left(k_{0}+p\right)d}\right)\left(e^{-i\phi}+e^{-i\left(k_{0}+2\bar{k}_{i}-p\right)d}\right)}{\left(p+i2\gamma\right)^{m+1}\left(p-2\bar{k}_{i}-i2\gamma\right)^{n+1}}$
$=$ $-\frac{ie^{i\bar{k}_{i}\left(X_{1}+X_{2}\right)}}{\bar{k}_{i}+i2\gamma}\sum_{mn}\frac{1}{m!}\left(\frac{-i\gamma\cos\phi}{\bar{k}_{i}+i2\gamma}e^{i\left(k_{0}+\bar{k}_{i}\right)d}\right)^{m+n}\{\left(e^{-i2\phi}+e^{-i2\left(k_{0}+\bar{k}_{i}\right)d}\right)e^{\left(i\bar{k}_{i}-2\gamma\right)X_{mn}}\sum_{l=0}^{n}\tilde{C}_{mnl}\left(k\right)\Theta\left(X_{mn}\right)$
$+e^{-i\left(k_{0}+\bar{k}_{i}\right)L-i\phi}\sum_{\alpha=\pm}e^{\left(i\bar{k}_{i}-2\gamma\right)X_{mn}^{\alpha}}\sum_{l=0}^{n}\tilde{C}_{mnl}^{\alpha}\left(k\right)\Theta\left(X_{mn}^{\alpha}\right)\}+\left(X_{1}\longleftrightarrow X_{2}\right),$
with
$\tilde{C}_{mnl}\left(k\right)=\frac{\left(m+n-l\right)!}{\left(n-l\right)!l!}\left[-i2\left(\bar{k}_{i}+i2\gamma\right)\left(\left(X_{1}-X_{2}\right)+\left(m-n\right)L\right)\right]^{l},$ (S33)
and
$\tilde{C}_{mnl}^{\pm}\left(k\right)=\frac{\left(m+n-l\right)!}{\left(n-l\right)!l!}\left[-i2\left(\bar{k}_{i}+i2\gamma\right)\left(\left(X_{1}-X_{2}\right)+\left(m-n\pm 1\right)L\right)\right]^{l}.$ (S34)
Here, the effective distances $X_{mn}$ and $X_{mn}^{\pm}$ are defined as
$X_{mn}=X_{1}-X_{2}+\left(m-n\right)L$ (S35)
and
$X_{mn}^{\pm}=X_{1}-X_{2}+\left(m-n\pm 1\right)L.$ (S36)
In the limit where $\left|X_{1}-X_{2}\right|$ goes to infinity, the contribution to the two-photon wavefunction from the background fluorescence vanishes, and the wave function
$w_{t}\left(X_{1},X_{2}\right)\rightarrow\frac{1}{2\pi}t\left(k_{1}\right)t\left(k_{2}\right)\left(e^{ik_{2}X_{1}}e^{ik_{1}X_{2}}+e^{ik_{1}X_{1}}e^{ik_{2}X_{2}}\right).$ (S37)
Supplementary Note 4: $N$ coupling points
With $N$ coupling points with the same coupling strength $V$ and diffirent phase $\phi_{n}$, the interaction between the giant cavity and the waveguide is $H_{1}=\int_{-\infty}^{+\infty}\tilde{V}_{k}\left(b^{{\dagger}}A_{+k}+\mathrm{H.c.}\right)\mathrm{d}k\mathrm{,}$where
$\tilde{V}_{k}=V\sqrt{2\sum_{m,n=0}^{N}e^{i\left(\phi_{m}-\phi_{n}\right)}\cos\left[\left(m-n\right)\left(k_{0}+k\right)d\right]}$ (S38)
and the coupled mode $A_{+k}$ and free mode $A_{-k}$ are defined as
$\left(\begin{array}[]{c}A_{+k}\\ A_{-k}\end{array}\right)=U_{k}\left(\begin{array}[]{c}a_{L-k}\\ a_{Rk}\end{array}\right),$ (S39)
with the unitary transformation
$U_{k}=\frac{V}{\tilde{V}_{k}}\left(\begin{array}[]{cc}\sum_{n=1}^{N}e^{i\phi_{n}}e^{-i\left(n-1\right)\left(k_{0}+k\right)d}&\sum_{n=1}^{N}e^{i\phi_{n}}e^{i\left(n-1\right)\left(k_{0}+k\right)d}\\ \sum_{n=1}^{N}e^{-i\phi_{n}}e^{-i\left(n-1\right)\left(k_{0}+k\right)d}&-\sum_{n=1}^{N}e^{-i\phi_{n}}e^{i\left(n-1\right)\left(k_{0}+k\right)d}\end{array}\right).$ (S40)
The self-energy is acquired as
$\Sigma\left(\omega\right)=-i\gamma\left[N+2\sum_{m>n}^{N}\cos\left(\phi_{m}-\phi_{n}\right)e^{i\left(m-n\right)\left(k_{0}+\omega\right)d}\right]$ (S41)
where $\gamma=2\pi V^{2}$. For photon scattering, one only needs to replace the self-energy in the $N=2$ case by $\Sigma\left(\omega\right)$ in Eq. (S41), as well as changing $\eta_{Rk}$ by
$\eta_{Rk}=\sum_{n=1}^{N}e^{i\phi_{n}}e^{i\left(n-1\right)\left(k_{0}+k\right)d}.$ (S42)
For example, the transmission amplitude $t\left(k_{i}\right)=\left\langle k_{f}^{R}\right|S\left|k_{i}^{R}\right\rangle$ is $t\left(k_{i}\right)=\delta\left(k_{i}-k_{f}\right)-i\gamma\left|\eta_{Rki}\right|^{2}G\left(k_{i}\right)\delta\left(k_{i}-k_{f}\right)$, where the Green’s function $G\left(\omega\right)=\frac{1}{\omega-\Sigma\left(\omega\right)}$. The S-matrix for two photons with right-moving incident momenta $k_{1}$and $k_{2}$, and right moving output momenta $p_{1}$ and $p_{2}$ is
$S_{p_{2}p_{1},k_{2}k_{1}}$ $=$ $\left[\delta\left(k_{1}-p_{1}\right)\delta\left(k_{2}-p_{2}\right)+\delta\left(k_{1}-p_{2}\right)\delta\left(k_{2}-p_{1}\right)\right]t\left(k_{1}\right)t\left(k_{2}\right)$ (S43)
$-i\frac{\gamma^{2}}{\pi}\eta_{Rp_{2}}^{\ast}\eta_{Rp_{1}}^{\ast}\eta_{Rk_{2}}\eta_{Rk_{1}}G\left(k_{1}\right)G\left(k_{2}\right)T_{S}G\left(p_{1}\right)G\left(p_{2}\right)\delta\left(k_{1}+k_{2}-p_{1}-p_{2}\right),$
where the $T_{s}$-matrix is
$T_{S}=\frac{U}{1-i\frac{U}{2\pi}\int\mathrm{d}\omega G\left(\omega\right)G\left(k_{1}+k_{2}-\omega\right)}.$ (S44)
The condition where the dynamics of the giant system is Markovian is equivalent to the condition when $\tilde{V}_{k}$ independent of $k$. It gives
$\phi_{1}=\pm\frac{\pi}{2}$ (S45)
and
$\sum_{n=1}^{N-1}\cos\left(\phi_{n}-\phi_{n+1}\right)=0,\sum_{n=1}^{N-2}\cos\left(\phi_{n}-\phi_{n+2}\right)=0,\sum_{n=1}^{N-3}\cos\left(\phi_{n}-\phi_{n+3}\right)=0,$... (S46)
For simplicity, we have set $\phi_{N}=0$. When these conditions can be fulfilled, the dynamics of the giant system can be exactlydepicted by the master equation
$\partial_{t}\rho=-i\left[H_{S},\rho\right]+N\gamma\left(2b\rho b^{\dagger}-\left\{b^{\dagger}b,\rho\right\}\right),$ (S47)
where $\gamma=2\pi V^{2}$. The input-output formalism read
$a_{out}^{l}\left(t\right)=a_{in}^{l}\left(t\right)-i\sqrt{\gamma}\sum_{n=1}^{N}e^{-i\phi_{n}}e^{i\left(n-1\right)k_{0}L}b\left(t-\left(n-1\right)d\right),$ (S48)
$a_{out}^{r}\left(t\right)=a_{in}^{r}\left(t\right)-i\sqrt{\gamma}\sum_{n=1}^{N}e^{-i\phi_{n}}e^{-i\left(n-1\right)k_{0}L}b\left(t+\left(n-1\right)d\right).$ (S49)
With $M=3$, the conditions can be fulfilled with $\phi_{1}=\pm\pi/2$, $\phi_{2}=\mp\pi/4$, resulting in $\tilde{V}_{k}=\sqrt{6}V$. In this case, the master equation reads
$\partial_{t}\rho=-i\left[H_{S},\rho\right]+3\gamma\left(2b\rho b^{\dagger}-\left\{b^{\dagger}b,\rho\right\}\right),$ (S50)
with the input-output formalism as
$a_{out}^{l}\left(t\right)=a_{in}^{l}\left(t\right)-i\sqrt{\gamma}\left[e^{-i\phi_{1}}b\left(t\right)+e^{-i\phi_{2}}e^{ik_{0}d}b\left(t-d\right)+e^{i2k_{0}L}b\left(t-2d\right)\right],$ (S51)
$a_{out}^{r}\left(t\right)=a_{in}^{r}\left(t\right)-i\sqrt{\gamma}\left[e^{-i\phi_{1}}b\left(t\right)+e^{-i\phi_{2}}e^{-ik_{0}d}b\left(t+d\right)+e^{-i2k_{0}L}b\left(t+2d\right)\right].$ (S52)
We note that for $N>2$, the chiral condition is not as simple as in the $N=2$. In fact, for a given incident momentum $k_{i}$, one need to solve a transcendental equation $\sum_{n=1}^{N}e^{i\phi_{n}}e^{-i\left(n-1\right)\left(k_{0}-k\right)d}=0$ to determine the phase $\phi_{n}$.